Solve p2 + 2py cotx = y2.
1
Expert's answer
2019-04-29T10:20:40-0400
"(y')^2+2y'ycotx-y^2=0"

"((y')^2+2y'ycotx-y^2)\/y^2=0"

"(y'\/y)^2+2y'cotx\/y-1=0"

"z=y'\/y"

"z^2+2zcotx-1=0"

"D=(2cotx)^2+4=4(1+(cotx)^2)=4\/(sinx)^2"

So we have two cases.


1)


"z=(-2cotx+\sqrt{4/(sinx)^2})/2=(1-cosx)/sinx"

"y'\/y=(1-cosx)\/sinx"

"dy\/y=(1-cosx)dx\/sinx"

"\intop dy/y=\intop (1-cosx)dx/sinx"

"lny=\intop (1-cosx)dx/sinx"

"\intop (1-cosx)dx/sinx=-\intop (1-cosx)d(cosx)/(sinx)^2="

"=-\intop (1-cosx)d(cosx)/(1-(cosx)^2)=-\intop d(cosx)/(1+cosx)="

"=-ln(1+cosx)+C"

So we get:



"lny=-ln(1+cosx)+C=ln(1\/(1+cosx))+C"

"y(x)=c\/(1+cosx)"

2)


"z=(-2cotx-\sqrt{4/(sinx)^2})/2=-(1+cosx)/sinx"

"lny=-\intop (1+cosx)dx/sinx=-ln(1-cosx)+C"

"lny=ln(1\/(1-cosx)+C"

"y(x)=c\/(1-cosx)"